# Importance of relativistic pericenter precession in identifying the presence of a third body near eccentric binaries

### v1

@@ IV.2 Maximum separation of the third body to which the perturbation effects can be detected
Figure 7 illustrates $|\Delta t|$ as a function of the number of GW bursts for different values of $R$ and $p_{0}$. Each color corresponds to a different value of $R$. Solid lines represent results for an initial value of $p_{0}=200m$, while dotted lines correspond to $p_{0}=100m$. The value of $\Delta t$ increases as $p_{0}$ increases. This behavior arises because the perturbative contributions in Eqs. (19), (20), (21), (II.3) scale as $p^{3}$, therefore for the fixed values of $m_{3}$ and $R$, a binary with a larger $p_{0}$ experiences a stronger perturbation. As a result, the third body’s influence is more pronounced in wider binaries. For the same $m_{3}$ and $R$, a wider inner binary accumulates a larger phase shift over time due to the stronger cumulative influence of the perturbation. Quantitatively, for $p_{0}=200m$, the timing effects induced by the third body can be detected out to distances of $R=5\times 10^{8}m$. For $p_{0}=100m$, the detectability range reduces to $R=10^{8}m$.
IV.3 The effect of binary’s center of mass motion on GW burst timing
In addition to the tidal force from the third body, binary’s COM around the barycenter of the triple system changes the distance between the source and the detector, altering the arrival time of the bursts in the detector frame. Generally Doppler-induced modulations are larger compared to tidal-induced modulations, however, Doppler-induced modulations can become suppressed depending on the location of the binary with respect to the detector. Under the assumption that the COM of the triple system is stationary, the observed timings of the GW bursts are given as [75]
@@ IV.3 The effect of binary’s center of mass motion on GW burst timing
where $t_{n}^{\rm int}$ is the intrinsic timing of bursts, $\bm{R}(t^{\rm int}_{i})$ is the separation vector between the inner binary and the third body. The vector $\bm{n}_{\rm obs}$ fixes the plane of the triple system with respect to the observer. It requires two angles: the inclination angle between the detector and the source plane $(\iota_{3})$ and the angle of the line of nodes $(\Lambda_{3})$. [^4] Note that the angle $\Lambda_{3}$ is degenerate with the other two planar angles $V_{3}$ and $\omega_{0}$ and can be dropped from Eq. (23) with an appropriate choice of coordinate system. However, we keep $\Lambda_{3}$ for better visualization. The vector $\bm{n}_{\rm obs}$ is given by
@@
Figure 8: The $\Delta t$ for intrinsic GW timing (blue) and the observed GW timing (red). We choose the same binary parameters as in Fig. 6 with $m_{3}=10^{6}m$ and $R=10m_{3}$. The observed timing can become suppressed depending on the location of the binary.
To study the effects of Roemer delay on the burst timing, consider a system similar to a stellar mass BBH orbiting a SMBH. We consider the same binary as in Fig. 6 with $m_{3}=10^{6}m$ and $R=10m$. Figure 8 illustrates $\Delta t$ corresponding to intrinsic GW timing and the observed timing. While the intrinsic GW timing is independent of location of the observer, the observed timing depends on the inclination angle $(i_{3})$ and the angle of line of nodes $\Lambda_{3}$. Note that the intrinsic $\Delta t$ (purely due to tidal effects) increases as a function of time. The observed $\Delta t$ (with Doppler effects) can become suppressed depending on the location of the binary. Therefore, depending on the location of the binary, tidal effects can be larger than the Doppler effects.
@@ V Summary and Conclusions
In this paper, we have studied the impact of conservative PN corrections on the dynamics of hierarchical triple systems consisting of a highly eccentric inner binary perturbed by a third body in a circular outer orbit. Highly eccentric binary emit GW radiation in the form of GW burst during each pericenter passage. The presence of a third body alters the timing of these GW bursts. We have incorporated 1PN corrections to the binary’s orbital evolution and the timing of GW bursts. The timing model developed in this study presents an accurate analytical description of the orbital evolution of such triple systems. Our results show that 1PN-induced precession significantly alters the binary’s orbital evolution.
In particular, our study demonstrates that the inclusion of 1PN corrections significantly modifies the timing of GW bursts in both perturbed and unperturbed binaries. The timing deviations due to 1PN corrections become very large compared to the typical measurement uncertainty on burst timing ($\sim 0.1$ sec for Einstein Telescope). Moreover, the inclusion of 1PN corrections leads to the rapid change of longitude of pericenter, that induce distinctive modulation features in the semilatus rectum and eccentricity of the inner binary, which arise only when the system is both perturbed and precessing. Such signatures represent potential indicators of a nearby third body.
We further showed that neglecting 1PN corrections in the timing model leads to significant systematic bias in the mass and distance of the third body. The errors in the quantity $m_{3}/R^{3}$ become as large as $\mathcal{O}(10^{-1}\mbox{--}10^{6})$, introducing substantial bias in the estimation of the tertiary’s properties. Additionally, in the absence of 1PN precession, the timing difference $\Delta t$ between perturbed and unperturbed binaries depends strongly on the initial argument of pericenter, leading to inaccurate predictions for certain values of $\omega_{0}$. However, when 1PN precession is included, this dependence becomes significantly weaker. This has important implications for detection: the GW timing signature induced by a third body becomes more stable and robust against variations in the initial orbital configuration.
@@
While our model captures key relativistic effects in system dynamics, we have made a few simplified assumptions. We have neglected the dissipative effects beyond 2.5PN order, which may introduce additional, potentially non-negligible contributions to the binary’s long-term evolution. We have also assumed that the outer orbit is circular and is coplanar with the inner binary, and we have omitted higher order PN corrections to the dynamics of the outer orbit. In a future study, it would be interesting to relax these assumptions and incorporate more realistic configurations – including eccentric or inclined outer orbits, and higher order PN corrections – to study their impact on system dynamics. Such extensions could reveal additional observational signatures [71, 72].
VI Acknowledgements
@@
Since the leading order effects of pericenter precession appear at 1PN order, the contribution of next-to-leading (2PN) order precession in the timing model is subdominant. Nevertheless, we provide the 2PN corrections to the timing model for completeness. The 2PN acceleration is given by [84]
@@
$B_{2}$ $=\frac{\dot{r}}{2}\Bigg{[}\eta\left(15+4\eta\right)v^{2}-\left(4+41\eta+8\eta^{2}\right)\frac{m}{r}-3\eta(3+2\eta)\dot{r}^{2}\Bigg{]}\,$, (26c)
The 2PN acceleration acts as a perturbing force and we calculate its components $\mathcal{R},\mathcal{S}$ and $\mathcal{W}$ in $\bm{n}$, $\bm{\lambda}$ and $\bm{e_{z}}$ basis, respectively, using Eq. (4) and are given as follows
@@
The secular change of orbital elements $p,e,w,t$ is obtained by integrating Eqs. (7) over a complete orbit. The 2PN corrections to $\omega$ and $t$ are given by the following equations
@@
Figure 9: The difference in timings due to 2PN corrections as a function of the number of bursts. The system parameters are the same as 3. For the perturbed binary, the difference between timings can reach up to $100$ sec at the $100^{\rm th}$ burst.

### v2

@@ IV.2 Maximum separation of the third body to which the perturbation effects can be detected
Figure 7 illustrates $|\Delta t|$ as a function of the number of GW bursts for different values of $R$ and $p_{0}$. Each color corresponds to a different value of $R$. Solid lines represent results for an initial value of $p_{0}=200m$, while dotted lines correspond to $p_{0}=100m$. The value of $\Delta t$ increases as $p_{0}$ increases. This behavior arises because the perturbative contributions in Eqs. (19) - (II.3) scale as $p^{3}$, and therefore for the fixed values of $m_{3}$ and $R$, a binary with a larger $p_{0}$ experiences a stronger perturbation. As a result, the third body’s influence is more pronounced in wider binaries. For the same $m_{3}$ and $R$, a wider inner binary accumulates a larger phase shift over time due to the stronger cumulative influence of the perturbation. Quantitatively, for $p_{0}=200m$, the timing effects induced by the third body can be detected out to distances of $R=5\times 10^{8}m$. For $p_{0}=100m$, the detectability range reduces to $R=10^{8}m$.
IV.3 The effect of the binary’s center of mass motion on GW burst timing
In addition to the tidal force from the third body, the binary’s COM around the barycenter of the triple system changes the distance between the source and the detector, altering the arrival time of the bursts in the detector frame. Generally Doppler-induced modulations are larger compared to tidal-induced modulations; however, Doppler-induced modulations can become suppressed depending on the location of the binary with respect to the detector. Under the assumption that the COM of the triple system is stationary, the observed timings of the GW bursts are given as [Isobel]
@@ IV.3 The effect of the binary’s center of mass motion on GW burst timing
where $t_{n}^{\rm int}$ is the intrinsic timing of bursts and $\bm{R}(t^{\rm int}_{i})$ is the separation vector between the inner binary and the third body. The vector $\bm{n}_{\rm obs}$ fixes the plane of the triple system with respect to the observer. It requires two angles: the inclination angle between the detector and the source plane $(\iota_{3})$ and the angle of the line of nodes $(\Lambda_{3})$. [^5] Note that the angle $\Lambda_{3}$ is degenerate with the other two planar angles $V_{3}$ and $\omega_{0}$ and can be dropped from Eq. (23) with an appropriate choice of coordinate system. However, we keep $\Lambda_{3}$ for better visualization. The vector $\bm{n}_{\rm obs}$ is given by
@@
Figure 8: The $\Delta t$ for intrinsic GW timing (blue) and observed GW timing (red). We choose the same binary parameters as in Fig. 6 with $m_{3}=10^{6}m$ and $R=10m_{3}$. The observed timing can become suppressed depending on the location of the binary.
To study the effects of Roemer delay on the burst timing, consider a system similar to a stellar mass BBH orbiting a SMBH. We consider the same binary as in Fig. 6 with $m_{3}=10^{6}m$ and $R=10m$. Figure 8 illustrates $\Delta t$ corresponding to the intrinsic GW timing and observed timing. While the intrinsic GW timing is independent of the location of the observer, the observed timing depends on the inclination angle $(\iota_{3})$ and the angle of line of nodes $\Lambda_{3}$. Note that the intrinsic $\Delta t$ (purely due to tidal effects) increases as a function of time. The observed $\Delta t$ (with Doppler effects) can become suppressed depending on the location of the binary. Therefore, depending on the location of the binary, the tidal effects can be larger than the Doppler effects.
@@ V Summary and Conclusions
In this paper, we have studied the impact of conservative PN corrections on the dynamics of hierarchical triple systems consisting of a highly eccentric inner binary perturbed by a third body in a circular outer orbit. Highly eccentric binaries emit GW radiation in the form of GW bursts during each pericenter passage. The presence of a third body alters the timing of these GW bursts. We have incorporated 1PN corrections to the binary’s orbital evolution and the timing of GW bursts. The timing model developed in this study presents an accurate analytical description of the orbital evolution of such triple systems. Our results show that 1PN-induced precession significantly alters the binary’s orbital evolution.
In particular, our study demonstrates that the inclusion of 1PN corrections significantly modifies the timing of GW bursts in both perturbed and unperturbed binaries. The timing deviations due to 1PN corrections become very large compared to the typical measurement uncertainty on burst timing ($\sim 0.1$ sec for Einstein Telescope). Moreover, the inclusion of 1PN corrections leads to the rapid change of the longitude of the pericenter, that induces distinctive modulation features in the semilatus rectum and eccentricity of the inner binary, which arise only when the system is both perturbed and precessing. Such signatures represent potential indicators of a nearby third body.
We further showed that neglecting 1PN corrections in the timing model leads to significant systematic bias in the mass and distance of the third body. The errors in the quantity $m_{3}/R^{3}$ become as large as $\mathcal{O}(10^{-1}\mbox{--}10^{6})$, introducing substantial bias in the estimation of the tertiary’s properties. Additionally, in the absence of 1PN precession, the timing difference $\Delta t$ between perturbed and unperturbed binaries depends strongly on the initial argument of the pericenter, leading to inaccurate predictions for certain values of $\omega_{0}$. However, when 1PN precession is included, this dependence becomes significantly weaker. This has important implications for detection: the GW timing signature induced by a third body becomes more stable and robust against variations in the initial orbital configuration.
@@
While our model captures key relativistic effects in system dynamics, we have made a few simplified assumptions. We have neglected the dissipative effects beyond 2.5PN order, which may introduce additional, potentially non-negligible contributions to the binary’s long-term evolution. We have also assumed that the outer orbit is circular and is coplanar with the inner binary, and we have omitted higher order PN corrections to the dynamics of the outer orbit. In a future study, it would be interesting to relax these assumptions and incorporate more realistic configurations – including eccentric or inclined outer orbits, and higher order PN corrections – to study their impact on system dynamics. Such extensions could reveal additional observational signatures [Hendriks:2024zbu, Hendriks:2024gpp].
Acknowledgements
@@
Since the leading order effects of pericenter precession appear at 1PN order, the contribution of next-to-leading (2PN) order precession in the timing model is subdominant. Nevertheless, we provide the 2PN corrections to the timing model for completeness. The 2PN acceleration is given by [PhysRevD.42.1123]
@@
$B_{2}$ $=\frac{\dot{r}}{2}\Bigg[\eta\left(15+4\eta\right)v^{2}-\left(4+41\eta+8\eta^{2}\right)\frac{m}{r}-3\eta(3+2\eta)\dot{r}^{2}\Bigg]\,.$ (26c)
The 2PN acceleration acts as a perturbing force, and we calculate its components $\mathcal{R},\mathcal{S}$ and $\mathcal{W}$ in $\bm{n}$, $\bm{\lambda}$ and $\bm{e_{z}}$ basis, respectively, using Eq. (4), which are given as follows:
@@
The secular change of orbital elements $p,e,w,t$ is obtained by integrating Eqs. (7) over a complete orbit. The 2PN corrections to $\omega$ and $t$ are given by the following equations:
@@
Figure 9: Difference in timings due to 2PN corrections as a function of the number of bursts. The system parameters are the same as those in Fig. 3. For the perturbed binary, the difference between timings can reach up to $100$ sec at the $100^{\rm th}$ burst.